OpenStudy (anonymous):

Write an epuation in slope-intercept form of the line that passes through the given points (3, -2) (6, 1)

OpenStudy (cwrw238):

slope m = 1-(-1) / (6-3) = 2/3 use standard equation y - y1 = m (x-x1) m=2/3, x1 = 6 and y1 = 1, so y - 1 = 2/3 (x-6) y = (2/3) x - 3 is the equation in slope /intercept form

OpenStudy (anonymous):

Your slope is wrong. Slope = (1-(-2))/(6-3) = 3/3 = 1.

OpenStudy (cwrw238):

oh sorry guys - i made an error the slope is 1 - (-2) / 6-3 = 1 y - y1 = m (x - x1) y - 1 = 1(x - 6) y = x - 5 is your equation

OpenStudy (anonymous):

y = x - 5 is correct.